# Target-driven Self-Distillation for Partial Observed Trajectories Forecasting

### v1

@@ 1 Introduction
For the autonomous driving system, predicting the trajectory of surrounding agents within the next few seconds is a crucial and challenging task. Accurate motion forecasting can enhance the safety of autonomous driving by promoting precise path planning and guiding the agents to follow safe routes, thereby rendering substantial support for ego-vehicle trajectory planning within the autonomous driving pipeline Jia et al. (2023); Hu et al. (2023); Wu et al. (2022), as well as aiding in robot navigation Huang et al. (2023); Chen et al. (2019) and tracking Cui et al. (2022).
Figure 1: Enhancing the robustness of the model to partially observed trajectory inputs through TSD. The agent requiring motion forecasting may have parts of historical trajectory occluded and thus unobservable. Through our proposed TSD, the model can adapt to this situation by leveraging predicted targets and self-distillation.
@@
Many carefully designed networks have achieved excellent results in motion forecasting, including recently proposed methods such as HiVT Zhou et al. (2022), QCNet Zhou et al. (2023), and SEPT Lan et al. (2023a), which have performed exceptionally well on large-scale motion forecasting datasets like Argoverse and Argoverse2. Despite their impressive performance, most open-source large-scale motion forecasting datasets contain carefully collected agent motion data, where the agents typically have complete input and output observations. However, in real-world situations, traffic safety is often compromised due to insufficient observations. For example, obstacles or other agents on the road can obscure parts of the agents, leading to incomplete observations when they appear, which in turn results in erroneous trajectory predictions and affects the safety of autonomous driving systems.
@@
Some methods Xu and Fu (2024); Sun et al. (2022) have designed specific networks to address the challenges of partial observation. However, if the goal is to retain the knowledge of the state-of-the-art network trained on fully observed data, many approaches adopt knowledge distillation Monti et al. (2022); Wang et al. (2024). Specifically, a network is first trained on fully observed data, and then the knowledge from this network is distilled into an identical network that receives only partially observed inputs. While the network obtained through knowledge distillation retains the original model’s structure and performs well under partial observation, this approach has two main drawbacks: first, the performance of the distilled network on fully observed data is lower than that of the originally trained network; second, the knowledge distillation process is cumbersome, requiring the training of one network followed by another distillation step, making the entire training process very expensive.
@@
To overcome the aforementioned challenges, as depicted in Figure 1 we propose a novel Target-driven Self-Distillation method (TSD) for motion forecasting. First, when observations are incomplete, if accurate target points can guide the trajectory prediction process, the model can still make accurate predictions despite the limited observations. Therefore, we propose an anchor-free target point generation method that accurately predicts the target points of agents, thereby guiding accurate trajectory prediction under both partial and full observations. Second, since partially observed trajectories can be viewed as a natural data-distortion branch, we refer to the design for classification tasks in Xu and Liu (2019) and develop a self-distillation method for motion forecasting. This method uses fully observed and partially observed inputs as two branches, jointly optimizes these branches, and employs empirical Maximum Mean Discrepancy (MMD) Long et al. (2015) as a non-parametric metric to measure the consistency of feature distributions between these distorted versions. Through these two aspects, we can obtain a robust model for partial observation data and outperform the vanilla version of the model under full observation with a single end-to-end training process.
@@
(1) We propose TSD, a novel plug-and-play method that enhances the robustness of motion forecasting models under partial observation trajectories, and simultaneously improves their motion forecasting capabilities under full observation trajectories.
(2) We desingn an anchor-free target point generation method based on the Transformer decoder, which iteratively predicts target points over both long-term and short-term time horizons. By accurately predicting target points, this method guides the model in performing precise trajectory prediction under partial observation conditions.
@@
(4) Extensive evaluations on multiple large-scale motion forecasting datasets demonstrate that our model not only significantly improves the robustness of baseline models in predicting motion under partial observation but also enhances prediction performance under full observation.
@@ 2 Related Works
2.1 Fully Observed Motion Forecasting
Fully Observed Motion Forecasting typically uses complete historical trajectories to predict future motion. Many methods have been proposed in current research, which can be categorized into anchor-based and anchor-free approaches. Methods like Chai et al. (2019),Zhao et al. (2021), and Afshar et al. (2024) cluster representative trajectory anchors as static anchors using unsupervised learning from training data. On the other hand, Varadarajan et al. (2022), Nayakanti et al. (2023), Shi et al. (2022), and Ngiam et al. (2022) generate dynamic, learnable anchors through implicit methods. After anchor generation, these methods combine the anchors with contextual information from the scene and use simple networks to predict future trajectories. Anchor-free methods, such as those using diffusion models like Jiang et al. , Li et al. (2023), and Gu et al. (2022), or generative networks like Barsoum et al. (2018), Kingma and Welling , and Gomez-Gonzalez et al. (2020), rely on the precise capturing of multimodal distributions for prediction. Additionally, some anchor-free methods adopt self-supervised learning approaches, such as Chen et al. (2023), Cheng et al. (2023) and Lan et al. (2023b). These methods learn more comprehensive scene interaction information by reconstructing features from masked regions and can also achieve impressive trajectory forecasting results. However, these methods, trained on carefully designed datasets with complete trajectory observations, exhibit poor robustness to partially observed trajectory inputs. Their performance typically degrades significantly when encountering those cases.
Our proposed plug-and-play method TSD can significantly enhance the robustness of these approaches when facing partially observed trajectories and also improves their performance under fully observation inputs.
@@ 2.1 Fully Observed Motion Forecasting
2.2 Partially Observed Motion Forecasting
Partially Observed Motion Forecasting refers to the use of partial historical data for future trajectory prediction. The previously mentioned Fully Observed Motion Forecasting methods tend to give better prediction results only on complete trajectories, but their performance is significantly degraded when partial historical trajectories are masked out, their performance decreases significantly. Numerous works have been proposed to address this issue. Some methods attempt to utilize knowledge distillation to mitigate this problem. Monti et al. (2022) and Wang et al. (2024) first train a teacher model on the full trajectory and use this model as the distillation target. Then, they ensure that the feature outputs of the student model, when the input is incomplete, match more closely with those of the teacher model under complete input conditions, thus enhancing performance in the case of trajectory data loss. Li et al. introduces a knowledge distillation framework for dynamically varying trajectory lengths and a dynamic neuron soft-masking mechanism to improve the model’s predictive performance when dealing with historical trajectory data of different lengths. Additionally, Xu and Fu (2024) learns representations from trajectory data of varying observation lengths and generates time-invariant representations, which are further optimized through an adaptive mechanism, thus improving prediction accuracy. Li et al. (2023) uses a bidirectional consistency diffusion model to predict future pedestrian trajectories based on short historical trajectory segments. Sun et al. (2022) proposes a unified feature extractor and a novel pre-training mechanism to capture relevant information in momentary observations, achieving pedestrian trajectory prediction even with extremely short observation lengths. However, these approaches often incur additional training costs and increased model complexity, and their performance remains poor under extreme missing data conditions.
In contrast, our proposed TSD leverages self-distillation to end-to-end training a model that is robust to partial observation trajectories, and is more efficient and cost-effective.
@@ 3 Problem Definition
Note that when it comes to partial motion forecasting, the initial observation of past state sequence ( $X_{\text{obs}}$ ) and the agent states ( A ) are not available at every frame; only a portion of the observations are fed into the model for motion forecasting.
@@ 4.1 Overview
The overview of our proposed framework is illustrated in Figure 2. We first employ an agent encoder to extract representations of the interested agent and surrounding agents, and a map encoder to extract representations of polylines and elements on the high-definition map. Subsequently, we use a sequential target point generator to iteratively predict long-term and short-term target points for the interested agent over different time horizons. These target points are then used to guide the trajectory predictor to iteratively predict multi-modal trajectories. Additionally, during the training phase, we add a partial observation branch. By generating masks, we create partially observed trajectories and feed them into the model along with fully observed trajectories for inference. The inference results from the partially observed trajectories are backpropagated in the same manner as those from fully observed trajectories, and the features extracted from partially observed trajectories are constrained by the features from fully observed trajectories through a loss function.
@@ 4.3 Target Generation and Trajectory Prediction
Here, we propose a target point generation method based on the Transformer Vaswani et al. (2017) decoder, which can iteratively generate sequential target points over both long-term and short-term time horizons.
Inspired by Zhou et al. (2023), we assume that the $x$ and $y$ coordinates of target points $\tau_{k}^{i}$ of the $k^{th}$ mode to be predicted, where $i\in\{1,2,\cdots,n\}$, follow a Laplace distribution, denoted as $\tau_{k,x}^{i},\tau_{k,y}^{i}\sim\mathcal{L}(\mu,b)$. Here, $n$ is the number of target points we set artificially for long-time horizon generation. Therefore, in our target point generation, the ultimate goal is to predict the parameters $\mu_{k,x}^{i},\mu_{k,y}^{i}$ and $b_{k,x}^{i},b_{k,y}^{i}$ for the $x$ and $y$ coordinates of each target point of every mode. Given the multimodal prediction of trajectories for $K$ modes, we generate $n$ target points for each mode; thus, the final set of target points generated for each agent is $\tau\in\mathbb{R}^{K\times n\times 4}$. For the dimension $n$, target points are generated sequentially over $n$ cycles, each time using the information gathered so far, rather than through an auto-regressive process. Eventually, we can concatenate them as a sequence.
@@
[…]
THOMAS Gilles et al. (2022) 2.16 0.88 1.51 0.20 1.95 4.71 0.64
GoRela Cui et al. (2023) 2.01 0.76 1.48 0.22 1.82 4.62 0.61
HPTR Zhang et al. (2023) 2.03 0.73 1.43 0.19 1.84 4.61 0.61
MTR++ Shi et al. (2023) 1.88 0.71 1.37 0.14 1.64 4.12 0.56
MacFormer Feng et al. (2023) 1.90 0.70 1.38 0.19 1.84 4.69 0.61
Gnet Gao et al. (2023) 1.90 0.69 1.34 0.18 1.72 4.40 0.59
Forecast-MAE Cheng et al. (2023) 1.91 0.69 1.34 0.17 1.66 4.15 0.59
ProphNet Wang et al. (2023b) 1.88 0.66 1.32 0.18 1.76 4.77 0.61
GANet Wang et al. (2023a) 1.81 0.62 1.19 0.14 1.64 4.09 0.55
QCNet Zhou et al. (2023) 1.78 0.62 1.19 0.14 1.56 3.96 0.55
TSD-Q (w/o ensemble) 1.87 0.63 1.25 0.15 1.68 4.33 0.58
TSD-Q (w/ ensemble) 1.78 0.61 1.17 0.14 1.59 4.10 0.56
[…]
dim. Method Obs. = 1 Obs. = 5 Obs. = 10 Obs. = 15 Obs. = 20
64 HiVT Zhou et al. (2022) 1.120/1.761/0.253 0.774/1.168/0.132 0.720/1.085/0.114 0.703/1.061/0.108 0.687/1.030/0.102
POP-H Wang et al. (2024) 1.228/1.857/0.227 0.742/1.117/0.118 0.704/1.060/0.107 0.692/1.040/0.104 0.693/1.043/0.103
TSD-H 1.040/1.600/0.211 0.739/1.098/0.117 0.694/1.031/0.104 0.679/1.009/0.098 0.675/0.998/0.096
128 HiVTZhou et al. (2022) 1.056/1.652/0.231 0.742/1.098/0.116 0.693/1.023/0.101 0.681/1.005/0.098 0.661/0.969/0.092
TSD-H 1.002/1.516/0.191 0.716/1.041/0.104 0.671/0.978/0.094 0.657/0.952/0.089 0.653/0.944/0.088
Table 2: Comparison the minADE6/minFDE6/MR6 performance of TSD-H, POP-H, and HiVT under partial observation trajectory inputs, conducted on the Argoverse motion forecasting dataset. For each hidden dimension, the best results are highlighted in bold.
@@ 4.4 Self-Distillation
(3) For the representation vectors of the intermediate features, we use Maximum Mean Discrepancy (MMD) Long et al. (2015) to reduce the differences between them. To match the feature distributions, a metric between the representation vectors of the two-branch distorted versions needs to be defined. We adopt the empirical MMD Long et al. (2015) as a non-parametric metric that has been widely used in domain adaptation to measure the discrepancy of distributions. We use the following empirical MMD loss to minimize the margin between the feature distribution obtained by full observation trajectories ($p(F_{A}),p(F_{X})$ and $p(F_{X\rightarrow A})$) with feature distribution obtained by partial observation trajectories ($p(F_{A}^{P}),p(F_{X}^{P})$ and $p(F_{X\rightarrow A}^{p})$) :
@@
where $n$ denotes the number of total features, $F_{A},F_{A}^{P}$ denote features of surrounded agents, $F_{X},F_{X}^{P}$ are features of the interested agent, and $F_{X\rightarrow A},F_{X\rightarrow A}^{P}$ are the interaction features between the interested agent and surrounded agents.
@@ 4.5 Training and Inference
We propose that the self-distillation process is only used during training. Specifically, during the training process, we adopt a winner-takes-all training strategy Lee et al. (2016). We select the trajectory with the best final target point prediction result from the complete observation for back-propagation and choose the trajectory with the same index from the partial observation output results. Overall, we denote loss for the full observation input branch as $\mathcal{L}_{ful}=\mathcal{L}_{tar}+\mathcal{L}_{reg}+\mathcal{L}_{cls}$, and the loss for the partial trajectory input branch is denoted as $\mathcal{L}_{par}=\mathcal{L}_{tar}+\mathcal{L}_{reg}+\mathcal{L}_{cls}$. Our overall training loss is as follows:

### v2

# Improving Partially Observed Trajectories Forecasting by Target-driven Self-Distillation

@@ 1 Introduction
For the autonomous driving system, predicting the trajectory of surrounding agents within the next few seconds is a crucial and challenging task. Accurate motion forecasting enhances safety by enabling precise path planning and guiding agents along safe routes, thereby strongly supporting ego-vehicle trajectory planning within the driving pipeline [jia2023think, hu2023planning, wu2022trajectory], as well as aiding in robot navigation [huang2023visual, chen2019crowd] and tracking [cui2022mixformer].
Figure 1: Motion forecasting models often encounter occluded historical trajectories. Our proposed TSD enhances robustness by leveraging self-distillation and predicting target to adapt to Partially Observation scenarios.
@@
Many carefully designed networks have achieved excellent results in motion forecasting, including recently proposed methods such as HiVT [hivt],DeMo [zhang2024decoupling], and LAformer [liu2024laformer], which have performed exceptionally well on large-scale motion forecasting datasets like Argoverse 1,Argoverse2 and Nuscenes. Despite their impressive performance, most open-source large-scale motion forecasting datasets contain carefully collected agent motion data, where the agents typically have complete input and output observations. However, in real-world deployments, the agent observation inputs received by trajectory prediction models typically originate from upstream objection tracking models. Due to interference from road obstacles and occlusions by dynamic vehicles, tracking results are highly prone to objection loss and localization errors, resulting in incomplete agent observation inputs. These inputs often exhibit random frame drops or partial trajectory absences, leading to erroneous trajectory predictions and consequently compromising the safety of autonomous driving systems.
@@
Some methods [xu2024adapting, sun2022human, qi2025robust, qi2025action] have designed specific networks to address the challenges of partial observation. However, if the goal is to retain the knowledge of the state-of-the-art network trained on fully observed data, many approaches adopt knowledge distillation [monti2022many, pop, wang2024adversarial, ye2025safedriverag]. Specifically, a network is first trained on fully observed data, and then the knowledge from this network is distilled into an identical network that receives only partially observed inputs. While the network obtained through knowledge distillation retains the original model’s structure and performs well under partial observation, this approach has two main drawbacks: first, the performance of the distilled network on fully observed data is lower than that of the originally trained network; second, the knowledge distillation process is cumbersome, requiring the training of one network followed by another distillation step, making the entire training process very expensive.
@@
To overcome the aforementioned challenges, as depicted in Figure 1 we propose a novel Target-driven Self-Distillation method (TSD) for motion forecasting. First, when observations are incomplete, if accurate target points can guide the trajectory prediction process, the model can still make accurate predictions despite the limited observations. Therefore, we propose an anchor-free target point generation method that accurately predicts the target points of agents, thereby guiding accurate trajectory prediction under both partial and full observations. Second, since partially observed trajectories can be viewed as a natural data-distortion branch, we refer to the design for classification tasks in [xu2019data, deng2025global, lv2025t2sg, zhu2023unsupervised] and develop a self-distillation method for motion forecasting. This method uses fully observed and partially observed inputs as two branches, jointly optimizes these branches, and employs empirical Maximum Mean Discrepancy (MMD) [mmd] as a non-parametric metric to measure the consistency of feature distributions between these distorted versions. Through these two aspects, we can obtain a robust model for partially observed data, and through an end-to-end training process, its performance is comparable to or even surpasses the original model under full observation conditions.
@@
(1) We propose TSD, a novel plug-and-play method which enhances the robustness of motion prediction models under partially observed trajectories while maintaining or even improving their motion prediction capabilities under fully observed trajectories.
(2) We design an anchor-free target point generation method based on the Transformer decoder. By accurately predicting target points, this method guides the model in performing precise trajectory prediction under partial observation conditions.
@@
(4) Experimental results on multiple large-scale datasets demonstrate that our approach not only significantly improves baseline robustness under partially observed scenarios, but also enhances the performance of some baselines in fully observed conditions, achieving comprehensive performance gains.
@@ 2 Related Works
Fully Observed Motion Forecasting Fully Observed Motion Forecasting typically uses complete historical trajectories to predict future motion. Many methods have been proposed in current research, which can be categorized into anchor-based and anchor-free approaches. Methods like [chai2019multipath],[zhao2021tnt], and [afshar2024pbp] cluster representative trajectory anchors as static anchors using unsupervised learning from training data. On the other hand, [varadarajan2022multipath++], [nayakanti2023wayformer], [shi2022motion], and [ngiam2022scenetransformerunifiedarchitecture] generate dynamic, learnable anchors through implicit methods. After anchor generation, these methods combine the anchors with contextual information from the scene and use simple networks to predict future trajectories. Anchor-free methods, such as those using diffusion models like [jiang2023motiondiffuser], [li2023bcdiff], and [gu2022stochastic], or generative networks like [barsoum2018hp], [kingma2013auto], and [gomez2020real], rely on the precise capturing of multimodal distributions for prediction. Additionally, some anchor-free methods adopt self-supervised learning approaches, such as [chen2023traj], [cheng2023forecast] and [lan2023sept]. These methods learn more comprehensive scene interaction information by reconstructing features from masked regions and can also achieve impressive trajectory forecasting results. However, these methods, trained on carefully designed datasets with complete trajectory observations, exhibit poor robustness to partially observed trajectory inputs. Their performance typically degrades significantly when encountering those cases. Our proposed plug-and-play method TSD can significantly enhance the robustness of these approaches when facing partially observed trajectories and maintain or even improve their performance under fully observed inputs.
@@
Partially Observed Motion Forecasting Partially Observed Motion Forecasting predicts future trajectories from Partially historical data. Fully Observed Motion Forecasting methods perform well on complete trajectories but their accuracy drops sharply when parts of the historical data are masked out.Numerous works have been proposed to address this issue. Some methods attempt to utilize knowledge distillation to mitigate this problem. [monti2022many] and [pop] use a teacher model trained with full trajectories to guide the student model, enhancing its prediction capability under Partially observation conditions. [li2024lakd] introduces a knowledge distillation framework for dynamically varying trajectory lengths and a dynamic neuron soft-masking mechanism to improve the model’s predictive performance when dealing with historical trajectory data of different lengths. Additionally, [xu2024adapting, qi2020stc] learns representations from trajectory data of varying observation lengths and generates time-invariant representations, which are further optimized through an adaptive mechanism, thus improving prediction accuracy. [li2023bcdiff] uses a bidirectional consistency diffusion model to predict future pedestrian trajectories based on short historical trajectory segments. [sun2022human, qi2021semantics] proposes a unified feature extractor and a novel pre-training mechanism to capture relevant information in momentary observations, achieving pedestrian trajectory prediction even with extremely short observation lengths. However, existing methods introduce additional training costs and only address scenarios with continuous missing input trajectories. In contrast, our proposed TSD method achieves end-to-end model training through self-distillation, significantly reducing computational overhead while effectively handling both continuous and random missing trajectory scenarios, demonstrating stronger robustness.
@@ 3 Problem Definition
Note that when it comes to partial motion forecasting, the initial observation of past state sequence ( $X_{\text{obs}}$ ) and the agent states ( A ) are not available at every frame;After our masking process, only a portion of the observed data is input into the motion prediction model. We categorize the partial trajectory masking of the target agent in the motion prediction task into two masking patterns: (1) Random Masking: This involves randomly selecting and masking certain frames in the complete historical sequence to simulate discrete trajectory gaps caused by objection loss during the operation of a objection tracking model. (2) Continuous Masking: This involves continuously masking the complete historical trajectory starting from the first frame to simulate scenarios where the target agent is occluded for an extended period, resulting in the absence of an entire trajectory segment.
@@ 4.1 Overview
The overview of our proposed framework is illustrated in Figure 2. We first use an encoder to extract representations of the target agent, surrounding agents, polylines, and elements on the high-definition map. Subsequently, we use a sequential target point generator to iteratively predict long-term and short-term target points for the interested agent over different time horizons. These target points are then used to guide the trajectory predictor to iteratively predict multi-modal trajectories. Additionally, during the training phase, we add a partial observation branch. By generating masks, we create partially observed trajectories and feed them into the model along with fully observed trajectories for inference. The inference results from the partially observed trajectories are backpropagated in the same manner as those from fully observed trajectories, and the features extracted from partially observed trajectories are constrained by the features from fully observed trajectories through a loss function.
@@ 4.3 Target Generation and Trajectory Prediction
Here, we propose a target point generation method based on the Transformer [vaswani2017attention] decoder, which can iteratively generate sequential target points over both long-term and short-term time horizons.
Inspired by [zhou2023query], we assume that the $x$ and $y$ coordinates of target points $\tau_{k}^{i}$ of the $k^{th}$ mode to be predicted, where $i\in\{1,2,\cdots,n\}$, follow a Laplace distribution, denoted as $\tau_{k,x}^{i},\tau_{k,y}^{i}\sim\mathcal{L}(\mu,b)$. Here, $n$ is the number of target points we set artificially for long-time horizon generation. Therefore, in our target point generation, the ultimate goal is to predict the parameters $\mu_{k,x}^{i},\mu_{k,y}^{i}$ and $b_{k,x}^{i},b_{k,y}^{i}$ for the $x$ and $y$ coordinates of each target point of every mode. Given the multimodal prediction of trajectories for $K$ modes, we generate $n$ target points for each mode; thus, the final set of target points generated for each agent is $\tau\in\mathbb{R}^{K\times n\times 4}$. For the dimension $n$, target points are generated sequentially over $n$ cycles, each time using the information gathered so far, rather than through an auto-regressive process. Eventually, we can concatenate them as a sequence.
@@ 4.4 Self-Distillation
(3) For the representation vectors of the intermediate features, we use Maximum Mean Discrepancy (MMD) [mmd] to reduce the differences between them. To match the feature distributions, a metric between the representation vectors of the two-branch distorted versions needs to be defined. We adopt the empirical MMD [mmd] as a non-parametric metric that has been widely used in domain adaptation to measure the discrepancy of distributions. We use the following empirical MMD loss to minimize the margin between the feature distribution obtained by full observation trajectories ($p(F_{A}),p(F_{X})$ and $p(F_{X\rightarrow A})$) with feature distribution obtained by partial observation trajectories ($p(F_{A}^{P}),p(F_{X}^{P})$ and $p(F_{X\rightarrow A}^{p})$) :
@@
where $n$ denotes the number of total features, $F_{A},F_{A}^{P}$ denote features of surrounded agents, $F_{X},F_{X}^{P}$ are features of the interested agent, and $F_{X\rightarrow A},F_{X\rightarrow A}^{P}$ are the interaction features between the interested agent and surrounded agents. MMD is adopted instead of KL divergence or L2 distance due to its theoretical and practical advantages in our self-distillation framework. As a kernel-based, non-parametric measure, MMD captures distributional discrepancies without assuming any parametric form, making it suitable for our task where the feature distributions of fully and partially observed trajectories are often non-Gaussian, multi-modal, and partially disjoint. In contrast, KL divergence is asymmetric and unstable when the distributions have non-overlapping support, while L2 distance overlooks the geometric structure of the feature space, leading to inferior alignment. Prior studies [huang2017like, xu2019data] have also verified MMD’s stability and effectiveness in knowledge distillation, confirming it as a theoretically grounded and empirically robust choice for aligning latent representations.
@@ 4.5 Training and Inference
We propose that the self-distillation process is only used during training. Specifically, during the training process, we adopt a winner-takes-all training strategy [lee2016stochastic, qi2019attentive]. We select the trajectory with the best final target point prediction result from the complete observation for back-propagation and choose the trajectory with the same index from the partial observation output results. Overall, we denote loss for the full observation input branch as $\mathcal{L}_{ful}=\mathcal{L}_{tar}+\mathcal{L}_{reg}+\mathcal{L}_{cls}$, and the loss for the partial trajectory input branch is denoted as $\mathcal{L}_{par}=\mathcal{L}_{tar}+\mathcal{L}_{reg}+\mathcal{L}_{cls}$. Our overall training loss is as follows:
@@
Dataset Method Fully Obs. Mask Rate. = 20% Mask Rate. = 40% Mask Rate. = 60% Mask Rate. = 80%
Argoverse 1 HiVT-64 0.687/1.030/0.102 0.735/1.113/0.120 0.793/1.207/0.138 0.859/1.331/0.167 0.950/1.489/0.200
HiVT-POP-64 0.693/1.043/0.103 0.732/1.103/0.121 0.790/1.199/0.137 0.842/1.283/0.155 0.894/1.361/0.171
HiVT-TSD-64 0.675/0.998/0.096 0.698/1.038/0.103 0.722/1.077/0.112 0.761/1.141/0.123 0.837/1.256/0.146
Argoverse 1 HiVT-128 0.661/0.969/0.092 0.713/1.059/0.107 0.783/1.188/0.135 0.860/1.330/0.162 0.931/1.454/0.189
HiVT-POP-128 0.672/0.989/0.096 0.713/1.066/0.109 0.760/1.144/0.123 0.812/1.233/0.144 0.873/1.334/0.165
HiVT-TSD-128 0.653/0.944/0.088 0.700/1.029/0.102 0.716/1.057/0.113 0.758/1.143/0.127 0.832/1.235/0.151
Dataset Method Fully Obs. Mask Rate. = 20% Mask Rate. = 40% Mask Rate. = 60% Mask Rate. = 80%
Argoverse 2 DeMo 0.680/1.354/0.155 1.199/2.132/0.268 1.976/3.352/0.372 1.999/3.146/0.379 2.197/3.429/0.416
DeMo-pop 0.685/1.349/0.156 0.913/1.648/0.218 1.527/2.694/0.359 2.447/4.217/0.460 2.399/4.032/0.449
Forecast-mae 0.712/1.408/0.217 1.488/2.729/0.354 2.703/4.957/0.511 3.719/6.761/0.579 3.655/6.529/0.562
DeMo-TSD 0.673/1.303/0.155 0.821/1.527/0.460 1.143/2.119/0.308 1.423/2.584/0.368 1.172/2.027/0.311
Dataset Method Fully Obs. Mask Rate. = 20% Mask Rate. = 40% Mask Rate. = 60% Mask Rate. = 80%
Nuscenes LAformer 1.197/2.292/0.365 1.356/2.422/0.379 1.565/2.611/0.405 1.740/2.793/0.424 2.001/3.163/0.456
LAformer-pop 1.235/2.350/0.376 1.248/2.376/0.382 1.267/2.405/0.386 1.298/2.542/0.391 1.421/2.589/0.409
LAformer-TSD 1.220/2.361/0.367 1.237/2.378/0.375 1.256/2.393/0.379 1.294/2.460/0.386 1.340/2.534/0.397
Table 1: We conducted comprehensive comparative experiments on three mainstream motion prediction benchmark datasets (Argoverse 1, Argoverse 2, and NuScenes) to evaluate the prediction performance of the original model and its enhanced versions using the POP and TSD methods under partially observed conditions with random masking. The experiments employed three core metrics—minADEk/minFDEk/MRk for evaluation, with k set to 6 for the Argoverse 1 and Argoverse 2 datasets, and k set to 5 for the NuScenes dataset. The best-performing results in all comparisons are highlighted in bold.
Dataset. Method Fully Obs. Obs. = 15 Obs. = 10 Obs. = 5 Obs. = 1
Argoverse 1 HiVT-64 0.687/1.030/0.102 0.774/1.168/0.132 0.720/1.085/0.114 0.703/1.061/0.108 1.120/1.761/0.253
HiVT-POP-64 0.693/1.043/0.103 0.692/1.040/0.104 0.704/1.060/0.107 0.742/1.117/0.118 1.228/1.857/0.227
HiVT-TSD-64 0.675/0.998/0.096 0.679/1.009/0.098 0.694/1.031/0.104 0.739/1.098/0.117 1.040/1.600/0.211
Argoverse 1 HiVT-128 0.661/0.969/0.092 0.681/1.005/0.098 0.693/1.023/0.101 0.742/1.098/0.116 1.056/1.652/0.231
HiVT-POP-128 0.672/0.990/0.096 0.675/1.017/0.101 0.688/1.040/0.106 0.742/1.105/0.119 1.062/1.675/0.237
HiVT-TSD-128 0.654/0.957/0.090 0.668/0.983/0.093 0.683/1.009/0.100 0.730/1.076/0.111 1.002/1.516/0.191
Table 2: Comparison the minADE6/minFDE6/MR6 performance of HiVT, HiVT-POP, and HiVT-TSD under Continuous Masking partial observation trajectory inputs, conducted on the Argoverse motion forecasting dataset. the best results are highlighted in bold.
5 Experiments
Figure 3: Qualitative results of HiVT and HiVT-TSD. The past trajectories are shown in yellow, the ground-truth trajectories are shown in red, and the predicted trajectories are shown in green. The white boxes are other vehicles around, and the red boxes are agents
5.1 Experiment Settings
Dataset. We evaluated our model on the Argoverse1 [argoverse1], Argoverse2 [wilson2023argoverse], and nuScenes [caesar2020nuscenes] datasets. Argoverse1 contains 323,557 driving scenarios (205,942/39,472/78,143 for train/val/test) with 5s sequences sampled at 10Hz. Models observe the first 2s and predict the next 3s. Argoverse2 provides 249,880 scenarios (199,908/24,988/24,984 for train/val/test) sampled at 10Hz, requiring 6s prediction based on the previous 5s. nuScenes includes 49,787 8s scenes (32,186/8,560/9,041 for train/val/test) at 2Hz, with 2s history to predict 6s future trajectories. All datasets provide agent trajectories and high-definition maps.
Metrics. We adopted evaluation metrics commonly used in previously state-of-the-art methods [zhou2023query, wang2023prophnet, feng2023macformer], i.e., minimum Average Displacement Error (minADEK), minimum Final Displacement Error (minFDEK), and Miss Rate (MRK). The minADE evaluates the $L2$ distance between the coordinates of each step of the top K best-predicted future trajectories by the model and the ground truth, averaged over all time steps. The minFDE evaluates the accuracy of the last point’s predicted coordinates. The MR metric evaluates the percentage of predictions in minFDE that exceed 2 meters.Only the top K predicted trajectories are evaluated if the model outputs more than K trajectories.For the Argoverse 1 and Argoverse 2 datasets, K is set to 6, while for the nuScenes dataset, K is set to 5.
Implement Details. We trained HiVT, DeMo, LAformer and Forecast-MAE with our method using PyTorch on 8 RTX 4090 GPUs, following the original hyperparameter settings.